OpenStudy (sabrina1):

Solve the following trigonometric equations on the interval [0; 2pi] : sin x + cos 2x = 1

myininaya (myininaya):

cos(2x)=cos^2(2x)-sin^2(x)=1-sin^2(x)-sin^2(x)=1-2sin^2(x)

myininaya (myininaya):

\[\sin(x)+1-2\sin^2(x)=1\]

myininaya (myininaya):

\[-2\sin^2(x)+\sin(x)=0\]

myininaya (myininaya):

\[\sin(x)(-2 \sin(x)+1)=0\]

myininaya (myininaya):

set both factors =0 and solve